OpenStudy (anonymous):

Vincent borrowed $500 from Mitchell for six months. How much interest will Mitchell earn if he charges Vincent a simple interest rate of 3 percent?

OpenStudy (mathmale):

the formula for simple interest is i=Prt, where i is the interest, P is the principal (amount borrowed or loaned), r is the interest rate (as a decimal fraction), and t is the time, in years.

OpenStudy (anonymous):

i =PRT/100 Therefore it is 500 x 6 x 3/100 which is 9000/100 which is 90 i guess the answer is 90 cuz a bit weak at it

OpenStudy (anonymous):

i=prt p=500, r=3%, t=6=12/6 so i=500*0.03*12/6=30

OpenStudy (anonymous):

stella83096 can you explain it

OpenStudy (anonymous):

explain about what point?

OpenStudy (anonymous):

how to got the answer to be as 30

OpenStudy (anonymous):

because simple interest formula is i=prt. and p is $500 and interest rate is 3% so its 0.03. and time is 6 months. but we should change into year.so divide by 12 its equal to 2. and times all of them

OpenStudy (anonymous):

K THANKSS GOT IT